Benzoic acid, C6H5COOH, is a weak acid with Ka = 6.3 × 10-5.

What amount of sodium benzoate, C6H5COONa, must be dissolved in 300. mL of 0.400 M C6H5COOH, in order to prepare a buffer of pH 4.50?

A. 0.24 mol

B. 0.12 mol

C. 0.36 mol

D. 0.60 mol

E. 0.48 mol
Answer #1

Number of moles of benzoic acid = M*V = 0.400 M * 0.300 L = 0.12 mol According to Henderson equation [base] pH = pka +log pH